OpenStudy (anonymous):

22. What is a simpler form of the following expression (-10x^3 + 30x - 20) / (-5x + 5)? (1 point) whats the answer? please i need this.

OpenStudy (anonymous):

\[ \frac{-10x^3+30x-20 }{ -5x+5 } \] \[ \frac{-10(x^3-3x+2) }{ -5x+5}\ \] \[\frac{ -10( \cancel{(x-1)}(x+2)(x-1) }{-5\cancel{(x-1)}} \] \[=2x^2+2x-4 \]